The radius of the sphere is 10 units.
What is the approximate volume of the sphere? Use π = 3.14 and round to the nearest whole cubic unit.

Answer by Boreal(15235)

V=4/3pi*r^3, r=10; r^3=1000
V=4/3pi*1000=4000pi/3=12,560 units^3/3=4187 to nearest cubic unit